Madison Company's cash ledger reports the following for the month ending March 31.

Date Amount No. Date Amount
Deposits: 3/4 $ 1,000 Checks: 541 3/2 $ 5,500
3/11 1,700 542 3/8 900
3/18 4,000 543 3/12 2,700
3/25 3,700 544 3/19 1,300
Cash receipts: 3/26-3/31 2,100 545 3/27 300
$ 12,500 546 3/28 600
547 3/30 1,600
Balance on March 1 $ 6,400 $ 12,900
Receipts 12,500
Disbursements (12,900 )
Balance on March 31 $ 6,000

Information from March's bank statement and company records reveals the following additional information:

  1. The ending cash balance recorded in the bank statement is $8,041.
  2. Cash receipts of $2,100 from 3/26–3/31 are outstanding.
  3. Checks 545 and 547 are outstanding.
  4. The deposit on 3/11 included a customer's check for $420 that did not clear the bank (NSF check).
  5. Check 543 was written for $3,300 for office supplies in March. The bank properly recorded the check for this amount.
  6. An automatic withdrawal for March rent was made on March 4 for $1,050.
  7. Madison's checking account earns interest based on the average daily balance. The amount of interest earned for March is $57.
  8. Last year, one of Madison's top executives borrowed $4,050 from Madison. On March 24, the executive paid $4,330 ($4,050 borrowed amount plus $280 interest) directly to the bank in payment for the borrowing.
  9. The bank charged the following service fees: $36 for NSF check, $15 for automatic withdrawal for rent payment, and $25 for collection of the loan amount from the executive.

pt.2A

Journal entry worksheet

  • Record the entries that increase cash.

Note: Enter debits before credits.

Transaction General Journal Debit Credit

pt.2 B

Journal entry worksheet

  • Record the entries that decrease cash.

Note: Enter debits before credits.

Transaction General Journal Debit Credit
Answer #1
2A
General Journal Debit Credit
Cash 4387
     Notes Receivable 4050
     Interest Revenue 337 =280+57
2B
General Journal Debit Credit
Accounts Receivable 420
Office Supplies 600 =3300-2700
Rent expense 1050
Service charge expense 76 =36+15+25
      Cash 2146
Alternatively, charge for NSF check may be added to Accounts Receivable
General Journal Debit Credit
Accounts Receivable 456 =420+36
Office Supplies 600
Rent expense 1050
Service charge expense 40 =15+25
      Cash 2146